# Taming Multimodal Joint Training for High-Quality Video-to-Audio Synthesis

### v1

@@ 4.3. Ablations
To elucidate the benefits of joint multimodal training, we mask away some modalities during training and observe the effects on the results.
The results are summarized in Section 4.2.
@@
This suggests that having a joint “text feature space” is beneficial for multimodality training.
@@
Note, we always sample audio-visual data and audio-text data at a 1:1 ratio except when we use no audio-text data (then we only train on audio-visual data).
When we do not use audio-text data, we observe overfitting and stop training early.
@@
1) our default of using the conditional synchronization module (Sec. 3.4);
@@
We illustrate this architecture in Section E.5;
@@ 4.4. Limitations
Our model generates unintelligible mumbles when prompted to generate human speech (from seeing a human talk or from text input).
@@ Appendix B Comparisons with Movie Gen Audio
Appendix C Details on Data Overlaps
@@
Appendix C summarizes the observed overlaps.
@@
We have carefully removed from our training data (AudioSet [8], AudioCaps [22], Freesound [31], WavCaps [43], and VGGSound [1]) anything that overlaps with any of the test sets (VGGSound and AudioCaps).
@@
Appendix D Details on the Audio Latents
@@
Appendix E Network Details
E.1. Model Variants
@@ E.3. Gating
E.4. Details on Synchronization Features
@@
E.5. Illustration of the “sum sync with visual” Ablation
@@
E.6. Visualization of Aligned RoPE
@@
Appendix F Training Details
We use a base learning rate of -, with a linear warmup schedule of 1K steps to train our model for 300K iterations.
We use the AdamW optimizer [24, 39] with and .
We notice occasional training collapse (to NaN) if the default was used instead.
We reduce the learning rate to - after 80% of the training steps, and once again to - after 90% of the training steps.
We use post-hoc EMA [20] with a relative width for all models.
For efficiency, we use bf16 mixed precision training in all training runs.
All the audio latents and visual embeddings are precomputed offline and loaded during training.
@@ Appendix F Training Details
Appendix G Additional Visualizations
@@
Below, we provide additional spectrogram visualizations comparing our method with prior works in Figures A6 and A7.

### v2

@@ 4.3. Ablations
To elucidate the benefits of joint multimodal training, we mask away some modalities during training and observe the effects on the results, summarized in Section 4.2.
@@
This suggests that having a joint “text feature space” is beneficial for multimodal training.
@@
We always sample audio-visual data and audio-text data at the same (roughly 1:1, see Appendix H) ratio except when we use no audio-text data. When we do not use audio-text data, we observe overfitting and stop training early.
@@
1) our default of using the conditional synchronization module (Section 3.4);
@@
We illustrate this architecture in Figure A4;
@@ 4.4. Limitations
Our model generates unintelligible mumbles when prompted to generate human speech (from seeing mouth movement or from text input).
@@ Acknowledgment.
We sincerely thank Kazuki Shimada and Zhi Zhong for their helpful feedback on this manuscript.
@@ Appendix B Comparisons with Movie Gen Audio
Appendix C Evaluation on the Greatest Hits Dataset
To address any potential bias by using the model-based DeSync metric, we conduct an additional experiment to assess temporal alignment by comparing the onsets of generated audio with ground-truth labels.
Concretely, we use the Greatest Hits [48] test set (244 videos) which contains videos with distinct and labeled sound events (a drumstick hitting <object>).
Notably, neither our models nor the baselines have been trained on this dataset.
We test both our method and baselines (using available code) on each video’s first 8 seconds (due to the models’ constraints):
we extract onsets from the generated audio following [6] and compare them with the labeled sound events.
We assess performance using accuracy, average precision (AP), and F1-score.
We provide the results in Section 4.2 and visualize the spectrograms in Figure A8.
MMAudio achieves significantly better performance in these model-free metrics.
Note that a high AP (not accuracy or F1) can be achieved by generating very few onsets (e.g., silent/noise), which is the case in Seeing&Hearing.
Appendix D Ablations on Filling in Missing Modalities
Among our training data, VGGSound is the only tri-modal (with class names as text) dataset while all others are audio-text.
For other data, we replace missing visual modalities (CLIP and Sync features) with end-to-end learnable embeddings ( and ) and missing text modalities with the empty string ().
We believe other methods to fill in missing modalities would be similarly effective since the deep net likely adapts.
Indeed, replacing the missing modalities with either all learnable embeddings or zeros yields no significant difference (Appendix D).
Note, we also drop modalities randomly during training to enable classifier-free guidance, which enhances the model’s robustness to missing modalities.
Ours 70.19 14.44 29.13 0.483
With all learnable 70.13 14.63 29.23 0.494
With zeros 69.91 14.60 29.22 0.496
Appendix E Details on Data Overlaps
@@
Appendix E summarizes the observed overlaps.
@@
We have carefully removed from our training data (AudioSet [9], AudioCaps [24], Clotho [5], Freesound [33], WavCaps [46], and VGGSound [1]) anything that overlaps with any of the test sets (VGGSound and AudioCaps).
@@
Appendix F Details on the Audio Latents
@@
Appendix G Network Details
G.1. Model Variants
@@ G.3. Gating
G.4. Details on Synchronization Features
@@
G.5. Illustration of the “sum sync with visual” Ablation
@@
G.6. Visualization of Aligned RoPE
@@
Appendix H Training Details
Training setup.
Unless otherwise specified, we used the same set of hyperparameters for all model sizes.
To train the models, we use the base learning rate of -, with a linear warm-up schedule of 1K steps, for 300K iterations, and with a batch size of 512.
We use the AdamW optimizer [26, 41] with , , and a weight decay of -.
If the default was used instead, we notice occasional training collapse (to NaN).
For learning rate scheduling, we reduce the learning rate to - after 80% of the training steps, and once again to - after 90% of the training steps.
For model exponential moving average (EMA), we use the post-hoc EMA [22] formulation with a relative width for all models.
For training efficiency, we use bf16 mixed precision training, and all the audio latents and visual embeddings are precomputed offline and loaded during training.
@@ Appendix H Training Details
Balancing multimodal training data.
Since we have significantly more audio-text training data (951K) than audio-text-visual data (180K), we balance the dataset by duplicating the audio-text-visual samples before random shuffling in each epoch.
By default, we apply a 5X duplication for a rough 1:1 data sampling ratio.
For the “medium” and “large” models, we reduce the duplication ratio to 3X to mitigate overfitting.
Duplicated videos.
We observe VGGSound dataset [1] contains duplicated videos, likely due to multiple uploads of the same video to YouTube under different video IDs.
For instance, videos 4PjEi5fFD6A (in training set) and FhaYvI1yrUM (in test set) are the same video.333Other uploads of this video that are not part of the VGGSound dataset include 1MQkMdlBezY and vHmRikW9axQ.
In Appendix E, we remove train-test sets overlaps by comparing the video IDs, though this method does not eliminate repeated uploads.
Since prior works have been trained on the same dataset, our training scheme remains a fair comparison.
Appendix I Additional Visualizations
@@
Below, we provide additional spectrogram visualizations comparing our method with prior works in Figures A6, A7 and A8.
@@
Figure A8:
Comparisons of prior works with MMAudio on the Greatest Hits [48] dataset.